A) What is the future value of this cash flow at 5​% interest rate at the end of year​ 7?

B) What is the future value of this cash flow at 9​% interest rate at the end of year​ 7?

C) What is the future value of this cash flow at 14​% interest rate at the end of year​ 7?
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

a.

A=13000*(1.05)^6+21000*(1.05)^5+29000*(1.05)^4+130,000

=$209,472.84(Approx).

b.

A=13000*(1.09)^6+21000*(1.09)^5+29000*(1.09)^4+130,000

=$225,049.27(Approx).

c.

A=13000*(1.14)^6+21000*(1.14)^5+29000*(1.14)^4+130,000

=$247,948.20(Approx).